OpenStudy (anonymous):

Which of the following is a polynomial with roots -sqrt 3 , sqrt 3, and -2

OpenStudy (anonymous):

x3 – 2x2 – 3x + 6 x3 + 2x2 – 3x – 6 x3 – 3x2 – 5x + 15 x3 + 3x2 – 5x – 15

OpenStudy (anonymous):

the one that you get when you multiply \[(x-\sqrt{3})(x+\sqrt{3})(x+2)\]

OpenStudy (anonymous):

\[(x-\sqrt{3})(x+\sqrt{3})(x+2)\] \[(x^2-3)(x+2)\] one more step and you are done

OpenStudy (anonymous):

is it answer b

OpenStudy (precal):

@MM6 multiply it out and we will check it for you